OpenStudy (anonymous):

find the domain of f(x)=(Sqrt x+3)/((x+8)(x-2))

OpenStudy (australopithecus):

Rules to memorize: -You cannot divide by zero -You cannot take the root of a negative number What values of x will violate these rules?

OpenStudy (anonymous):

well I think the answer I have hereis x>-3, x is not 2 is this correct?

OpenStudy (australopithecus):

Hint: You can take the root of zero

OpenStudy (anonymous):

x is greater or equal to 0?

OpenStudy (australopithecus):

\[\sqrt{0} = 0\]

OpenStudy (anonymous):

all real numbers than!

OpenStudy (australopithecus):

No, stop and think, you were close to the right answer before

OpenStudy (anonymous):

I came up with x is greater or equal to 0 is this right?

OpenStudy (australopithecus):

Well test your answer, sub +2 into your equation

OpenStudy (anonymous):

i did and got x is greater or equal to 0

OpenStudy (australopithecus):

The domain of a function is simply the values of x you can sub into an equation and acquire a real number as y If you subbed +2 as x into your function you will violate a rule I listed in my first post

OpenStudy (anonymous):

yea but i need to answer to verify please

OpenStudy (australopithecus):

\[f(2) = \frac{\sqrt{2+3}}{(2+8)(2-2)}\]

OpenStudy (australopithecus):

can you simplify that?

OpenStudy (anonymous):

yea and the answer to prove my theory is..?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

sorry 5/0

OpenStudy (australopithecus):

its not really a theory but I digress

OpenStudy (anonymous):

ok well what would be the final answer?

OpenStudy (anonymous):

it would be something hashtag 2 right?

OpenStudy (australopithecus):

I'm not here to give you the answer, I'm here to teach so I wont have to give you the answer

OpenStudy (australopithecus):

well if you found that 2 a possible value of x what does that tell you about the domain of f(x)

OpenStudy (anonymous):

all real numbers except -8,3 and 2!

OpenStudy (anonymous):

this is it right? just to make sure

OpenStudy (anonymous):

is this right?

OpenStudy (australopithecus):

Look at my rules again -You cannot divide by zero -You cannot take the root of a negative number Both of these rules apply directly to your problem, I recommend looking at each part of the expression and think about what x values will violate the rules I listed above. Then you will know what is in your domain and what isnt

OpenStudy (australopithecus):

I guess you can split this problem into two parts Look at the denominator (x + 8)(x - 2) you know this cannot equal zero otherwise you would divide by zero which is not allowed (if you are interested why just search on youtube same for the other rule). So, Set (x + 8)(x - 2) = 0 and solve for x Next look at the numerator \[\sqrt{x+3}\] we have a radical, and we know we cannot take the root of a negative number. So what x values will cause us to have a negative number under the radical?

OpenStudy (australopithecus):

Follow?